a parent who has not had to use this math since high school and my son came to me for help, Help!!! heres two types of problems he has to do 3(6+4) = 3(___)+ 3(4) and my son has to do some simplifys 12(x + 2) or 9 (p - 6)
Answer by wuwei96815(245)

3(6+4) = 3(__) + 3(4)

Rewrite the equation and solve:
3(6+4) = 3(x) + 3(4)
3(10) = 3x + 12
30 -12 = 3x
18/3 = x
6 = x

The second part of your question is unclear. Was it copied correctly?